C97C1A

PI: Sanghera
Title: CAL - DA193

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: A (dual polarisation)
Obs time (UT): EU 18:00(12/02)-24:00(12/02)
Stations: CmEbJbMcNtOnTrShUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M. Bondi at Wed Feb 12 19:32:18 EST 1997) Success
experiment appears to have ended succesfully.

Torun (Roman Feiler at Wed Feb 12 22:13:09 EST 1997) Success
No known problems

Medicina (T. Venturi at Thu Feb 13 05:45:18 EST 1997) Minor failures
We have just realised that VC13 was badly patched, i.e. on IF2 insted of IF1

Effelsberg (dgraham at Thu Feb 13 07:07:39 EST 1997) Failure
Lost 1930 to 2230 due to operator typing mistake

Onsala (John Conway at Thu Feb 13 07:18:33 EST 1997) Minor failures
Formatter may have jumped 100ns during obervations, jump occurred sometime in 24hrs after 1200UT on Feb 12th, investigating. Otherwise no known problems.

Cambridge (Alastair Gunn at Tue Feb 18 06:47:17 EST 1997) Success
No known problems. Log file says JBLOVELL but should be CAMBG32M. Errors occurring during Tsys measurements (investigating). Weather was overcast but calm and dry. Tape EVN00182,FA90 sent to Bonn via Woodcroft on 18/02/97.

Westerbork (Tony Foley at Tue Feb 18 09:41:52 EST 1997) Minor failures
Array had lots of restarts about 2300and 2345UT because of some online software problem

MPIfR (dag+wa at Tue Feb 18 10:30:04 EST 1997) Success
Fringes found on Effelsberg - Torun with expected delay from GPS, also Effelsberg formatter 1 sec slow. (this offset stays until 16th).

MPIfR (dgraham at Wed Feb 19 10:11:39 EST 1997) Minor failures
correlator News - fringes Eff-Jodrell-Cambridge but Jodrell crossed plarisation. Station notified

MPIfR (W. Alef \& D. Graham at Thu Feb 20 03:31:57 EST 1997) Minor failures
Fringes found for Effelsberg, Torun; Good recording quality for both. Torun has its 8 BBCs assigend to first 4 frequencies of both pola- rizations. Track 1 no data, BBC7 defect. Fringes found for Cambridge in expected upper 4 frequencies, thus hardly any overlap with Torun! Fringes found for Jodrell Bank, but polarizations crossed. Jodrell informed on Wednesday afternoon. Track 18 defect.

Jodrell (Alastair Gunn at Fri Feb 21 03:19:34 EST 1997) Minor failures
POLARISATIONS ARE CROSSED. VC01 returned bad numbers on Tsys measurements on this and all subsequent experiments. Weather was overcast but calm and dry. Tape EVN00033,72CD sent to Bonn via Woodcroft on 17/02/97.

MPIfR (dgraham at Fri Feb 21 06:28:48 EST 1997) Success
correlator news- fringes found WSRT, with phase problems, signal goes off and on at 10sec intervals and has randow phase jumps. Onsala found at predicted GPS value.

Shanghai (Huang Xinyong at Wed Mar 5 04:15:30 EST 1997) Success
Clock offset=-0.33 msec. No problems, weather was fine.

Urumqi (Lee Suqin at Tue Mar 18 13:59:47 MET 1997) Minor failures
The parity errors are high as data was recorded with the old head.

Calibration (Fredrik T Rantakyro at Thu May 8 11:25:01 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name c97c1a.antab